OpenStudy (anonymous):

Solve: −7(x − 2) = 42 −4 4 −6.3 −8

OpenStudy (anonymous):

−7(x − 2) = 42 -7x+14=42 multiply in brackets -7x+14-14=42-14 subtract 14 from each side to isolate the -7x -7x=28 -7x/-7=28/-7 divide by negative 7 to isolate x by it self and remove the negative on the left x= - 4

OpenStudy (anonymous):

Can I Get A Medal??

OpenStudy (anonymous):

Lol! :(

OpenStudy (anonymous):

Thanks!